• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Sparveris

    Log entry time 08:03:52 on May 5,2007

    Entry number 201031

    keyword=shift summary

    Spent 2 hours trying to get acceptable beam.

    Run quasielastic to check detectors.

    Around 02:30 we went into production mode (3He @ Q2=61fm-2 with 100uA beam) and have been running smoothly since.

    Day shift takes over.